If Belson Electronic's new converter produces less than the 110 volts of their old product, they must redesign it. A sample of 303 converters generated a standard deviation of 47.3 volts. Set Alpha at 5 percent. a mean of 107.2 volts, with State and test the proper set of hypotheses. Is a redesign necessary?
Answer by Boreal(15235)

Assume random sample and independence of process
Ho: voltage is 110
Ha: it is not 110
alpha=0.05 P{reject Ho|Ho true)
test is a t 0.975 df=302* s/sqrt(n)=(mu new-mu old)/s/sqrt(n)
critical value |t|>1.968
t=(107.2-110)/47.3/sqrt(303)
=-2.8*sqrt(303)/47.3
=-1.03
p=0.30
Fail to rejected Ho and insufficient evidence to redesign.